# Parallel Batch Scheduling With Incompatible Job Families Via Constraint Programming

### v2

@@ VI The Proposed CP Models
This section presents the three proposed CP models that directly handle the interruptions in their formulation. These models make explicit use of the ordered set of possible batches =$\mathcal{B}=\set{1,2,3\ldots}$. The family $f^{b}\in\mathcal{F}$ of each possible batch $b\in\mathcal{B}$ is defined a priori. In addition, the model only allows jobs $\mathcal{J}_{f^{b}}$ of family $f^{b}$ to be processed on batch $b$. Hence, this set can be further partitioned into the mutually exclusive sets $\mathcal{B}_{f}=\set{b\in\mathcal{B}:f^{b}=f}\subset\mathcal{B}$ of possible batches for each family $f\in\mathcal{F}$. In this way, the first $|\mathcal{B}_{1}|$ elements of $\mathcal{B}$ correspond to the possible batches where jobs of family $1$ can be processed; the next $|\mathcal{B}_{2}|$ elements to the possible batches for jobs of the family $2$, and so on. The number of possible batches of each family $|\mathcal{B}_{f}|$ is upper bounded by the number of jobs that belong to the family $|\mathcal{J}_{f}|$. On the other hand, assuming $l_{f}>0$, this number is also upper bounded by $\lceil\frac{N_{f}}{l_{f}}\rceil$, where $N_{f}=\sum_{j\in\mathcal{J}_{f}}s_{j}$ is the aggregated size of all the jobs of family $f$. Consequently, we set $|\mathcal{B}_{f}|=\min\set{}{\mathcal{J}_{f}|,\lceil\frac{N_{f}}{l_{f}}\rceil}$.
VI-A The Assign-and-Schedule Model
Model 2 The Assign-and-Schedule Model Inspired by [8]
@@ VI-A The Assign-and-Schedule Model
$y_{b}$$\in\{[a,a+p^{f^{b}}):a\in\mathbb{Z}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(10)$
$y_{bm}$$\in\{[a,a+p^{f^{b}}):a\in\mathbb{Z}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(11)$
@@
$\sum_{j\in\mathcal{J}_{f^{b}}}s_{j}\cdot z_{jb}\nobreak\leavevmode$ $\geq l_{f^{b}}\cdot\texttt{presenceOf}\left(y_{b}\right);$ $(17)$
$\sum_{j\in\mathcal{J}_{f^{b}}}s_{j}\cdot z_{jb}\nobreak\leavevmode$ $\leq u_{f^{b}}\cdot\texttt{presenceOf}\left(y_{b}\right);$ $(18)$
@@
Model 2 presents the Assign-and-Schedule (AS) model. It is based on the CP model by Nascimento et al. [8] for AM processes, which deals with a single compatible family and does not include capacity constraints. This paper extends their formulation to handle multiple incompatible families and include the necessary constraints to ensure the batch sizes. The model name originates from its clear distinction between assignment and scheduling decisions. While the job assignment to batches is handled with binary variables, the batch sequencing is handled with CP scheduling logic.
Equation (2) defines a binary variable $z_{jb}$ that takes the value of 1 if job $j\in\mathcal{J}$ is assigned to batch $b\in\mathcal{B}_{f_{j}}$, or 0 otherwise. Equation (2) defines an optional interval variable $y_{b}$ of size $p^{f^{b}}$ that represents batch $b\in\mathcal{B}$. It is optional because a batch with no jobs assigned should not be present in the solution. Equation (2) defines an optional interval variable $y_{bm}$, also of size $p^{f^{b}}$, that represents batch $b$ being processed on machine $m\in\mathcal{M}$. Equation (2) defines an interval sequence variable $\varphi_{m}$, which represents the sequence of batches scheduled on machine $m$. This sequence variable can be any permutation of the batch intervals $\set{y_{bm}}_{b\in\mathcal{B}}$ on the machine $m$ present in the solution.
Objective function (2) minimizes the TWCT. Constraints (2) assign each job to exactly one possible batch. Constraints (2) deactivate the presence of the batches if no jobs are assigned to them. Constraints (2) activate the presence of the batches if at least one job is assigned to them. Constraints (2)-(2) ensure that if a batch is used, its size is within the specified limits of the batch family. Constraints (2) guarantee that batches are scheduled once all the jobs in them are available. Constraints (2) schedules each batch on exactly one machine, if present. Finally, constraints (2) use the global constraint noOverlap, which ensures that the present interval variables in the provided sequence do not overlap. Hence, constraints (2) ensure that only one batch is processed at a time on each machine.
Figure 2 presents the optimal solution of the AS model that has a TWCT of $1,700$. The increase in the TWCT from the optimal (but infeasible) solution of the NS model is the result of a more restrictive model that enforces correctly assigned jobs to no overlapping batches. In the optimal solution of the problem, Jobs 1 and 3 are assigned to Batch 1, while Jobs 2 and 4 are assigned to Batch 2, and the batch interval variables do not overlap on the machine.
@@
Although the beauty of the AS model is its clear distinction between assignment decisions with binary variables and scheduling decisions with interval variables and global constraints, this separation reduces the ability of the CP solver to prune the domains of the scheduling variables. The noOverlap constraints, in charge of handling the scheduling logic, provide the CP engine with a global view of the disjunctive constraints and effectively reduce the search space. However, constraints (2)-(2), in charge of handling the assignment while respecting the batch limits, are local in nature, reduce the overall propagation, and provide weak information to the global scheduling constraints. To remedy this, the models presented next make heavy use of global constraints for both the assignment and scheduling components.
VI-B The Synchronized Model
Model 3 The Synchronized Model Inspired by [10]
@@ VI-B The Synchronized Model
$y_{bm}$$\in\{[a,a+p^{f^{b}}):a\in\mathbb{Z}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(23)$
@@
${\mathscr{l}}_{m}^{b}$ $=\sum_{j\in\mathcal{J}_{f^{b}}}\text{pulse}\left(x^{b}_{jm},s_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},b\in\mathcal{B}.$ $(25)$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}}:$ (32)
alwaysIn $\bigl{(}{\mathscr{l}}_{m}^{b},x_{jm}^{b},l_{f^{b}},u_{f^{b}}\bigr{)}.$
Model 3 presents the Synchronized (S) model. It is based on the CP model by Cakici et al. [10] for AM processes, which deals with a single compatible family; this paper extends their formulation to handle multiple incompatible families. Equation (3) defines an optional interval variable $x_{jm}^{b}$, of size $p_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m\in\mathcal{M}$. Equation (3) defines an optional interval variable $y_{bm}$ of size $p^{f^{b}}$ that represents the option of batch $b\in\mathcal{B}$ being processed on machine $m$. Equation (3) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (3) defines a cumulative function ${\mathscr{l}}_{m}^{b}$ that represents the load on machine $m$ due to the jobs in batch $b$.
Objective function (3) minimizes the TWCT. Constraints (3) assign each job to exactly one batch on exactly one machine. Constraints (3) allow each batch to be processed in at most one machine. Constraints (3) use the synchronize global constraint, which takes two inputs: an interval variable $v=x_{jm}^{b}$ and a set of interval variables $B=\set{y_{bm}}$. If $v$ is present, this constraint attempts to synchronize it with any intervals in $B$ that are also present [38], i.e., $\texttt{presenceOf}(v)=1\implies\texttt{startOf}(v)=\texttt{startOf}(b)\wedge\texttt{endOf}(v)=\texttt{endOf}(b),\forall\leavevmode\nobreak\ b\in B$ s.t. $\texttt{presenceOf}(b)=1$. In practice, when the interval variable representing the job being assigned to a batch on a machine (i.e., $x_{jm}^{b}$) is part of the solution, the synchronize constraint attempts to align its start and end times with those of the batch on the machine (i.e., $y_{bm}$), if it is also present. However, the synchronize constraint only ensures synchronization if both interval variables are present; it does not inherently enforce the presence of $y_{bm}$ when $x_{jm}^{b}$ is scheduled. This subtle issue, where the presence of $x_{jm}^{b}$ should logically imply the presence of $y_{bm}$ is addressed in the original model by Cakici et al. [10] by means of an additional constraints that defines the correct duration of the batches depending on the jobs assigned to it, if any. This constraint is necessary in their AM application as the duration of the batch depends on the size of the parts 3D-printed in it. However, this constraint is not required in the context of the diffusion area studied in this paper since the duration of the batches is already predefined. Nonetheless, the presence of the correct batches is still needed to be enforced. To resolve this, constraints (3) introduce an explicit implication constraint which ensures that, whenever $x_{jm}^{b}$ is included in the solution, the corresponding batch interval $y_{bm}$ is also present. Constraints (3) use the noOverlap global constraint to ensure that the present interval variables in the machine sequence do not overlap, guaranteeing at most one batch being processed at a time on each machine. Finally, constraints (3) guarantee that the load of the machines when processing each job in its batch is within the size limits of the batch’s family.
Figure 3 shows the optimal solution of the S model, which has the same TWCT of $1,700$ as the AS model. Note how the S model enforces all the jobs in the same batch to start and end together in a synchronized fashion. In the context of the diffusion area in a wafer fab, this solution allows the thermal stability required by the finished products.
@@
The S model inspired by Cakici et al. [10] replaces the binary variables of the AS model with interval variables and uses more global constraints, like the synchronize and alwaysIn constraints. These changes make the S model more efficient than the AS model, as evidenced in the computational experiments. Nonetheless, some constraints of the S model still handle the assignment of jobs to batches through the binary presence variables of those intervals. This lack of global constraints to deal with the assignment in the model by Cakici et al. [10] leads to a weak propagation between the assignment and scheduling parts of the model. The novel CP model presented next exclusively uses global constraints to enhance computational performance.
VI-C The Redundant Synchronized Model
Model 4 presents the Redundant Synchronized (RS) model. It exclusively uses global constraints that exploit the problem structure to better guide the search process. Furthermore, it includes redundant variables and global constraints that increase the computational performance.
Equation (4) defines an interval variable $x_{j}$ of size $p_{j}$ that represents job $j\in\mathcal{J}$. Equation (4) defines an optional interval variable $x_{jm}$ that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. Equation (4) defines an optional interval variable $x_{jm}^{b}$, also of size $p_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m$. Equation (4) defines an optional interval variable $y_{b}$ of size $p^{f^{b}}$ that represents batch $b\in\mathcal{B}$. Equation (4) defines an optional interval variable $y_{bm}$ of size $p^{f^{b}}$ that represents the option of batch $b$ being processed on machine $m$. Equation (4) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (4) defines the set $V_{jm}^{b}$ of interval variables that have to be synchronized with interval $x_{jm}^{b}$, if it is present. Equation (4) defines the unique state ${\mathfrak{f}}_{m}$ of the machine $m$ at every point in time. Equation (4) defines a cumulative function ${\mathscr{l}}_{mf}$ that represents the load on machine $m$ with jobs of family $f$.
Model 4 Redundant Synchronized model

### v3

@@ VI The Proposed CP Models
This section presents the four proposed CP models that can be implemented in other commercial and open-source solvers.
VI-A The Automaton model
The Automaton (AU) model derives its name from the automaton global constraint it relies on, which is available in different open-source CP solvers [41, 22, 42]. Automaton constraints model finite state machines, enabling the representation of feasible transitions between states. In the AU model, these finite states represent the elapsed time when processing a batch of each family. Consider the automaton depicted in Figure 2, where nodes are states and arcs are transition possibilities. The descriptions in each state help understand the figure and the numbers in parenthesis ultimately identify them. Each arc is identified by a triplet $(i,k,\tau)$, where $i$ is the initial state, $k$ is the end state, and $\tau$ is the required transition value to trigger such transition from $i$ to $k$. For example, if a machine is empty and receives a transition value of 0, then the machine remains empty. But if it receives a transition value of 1, it means that a new batch of family 1 starts on the machine. From this point, the only possible transition values that the automaton can receive are the necessary to progress the states until completing the processing time ${\delta}^{1}$ of the batch. Once the batch is completed, three possible transitions can be made: either the machine becomes empty (0), or a new batch of family 1 starts (1), or a new batch of family 2 starts (${\delta}^{1}+1$). From that point, a new sequence of transition values must be given such that the state transitions respect the automaton structure.
Figure 2: Example automaton with two job families
Let $\mathcal{G}=(\mathcal{N},\mathcal{A})$ be the graph describing the automaton. Let $\mathcal{N}=\set{i:0\leq i\leq\sum_{f\in\mathcal{F}}{\delta}^{f}}$ be the nodes that represent finite states, where 00 represents the empty state. Note that each family has a sequence of states. Let $\sigma_{f}\in\mathcal{N}$ and $\varphi_{f}\in\mathcal{N}$ be the initial and end states in the sequence of states of family $f\in\mathcal{F}$, respectively. By setting ${\delta}^{0}=0$, then $\sigma_{f}={\delta}^{f-1}+1$ and $\varphi_{f}={\delta}^{f-1}+{\delta}^{f}$. In this way, let $\Lambda=\set{\sigma_{f}:f\in\mathcal{F}}$ and $\Omega=\set{\varphi_{f}:f\in\mathcal{F}}$ be the sets of all initial and final states of the families, respectively. Let $\mathcal{A}=\set{(i,k,\tau):i,k\in\mathcal{N},\tau\in\mathbb{Z}^{1}_{+}}$ be the set of arcs that identify the possible transitions between states and the transition values that trigger them. This set of transition triplets can be further divided in five subsets, i.e., $\mathcal{A}=\mathcal{A}_{0}\cup\mathcal{A}_{1}\cup\mathcal{A}_{2}\cup\mathcal{A}_{3}\cup\mathcal{A}_{4}$. Let $\mathcal{A}_{0}=\set{(0,0,0)}$ be the transition triplet that allows an empty machine remain empty. Let $\mathcal{A}_{1}=\set{(0,\sigma_{f},\sigma_{f}):f\in\mathcal{F}}$ be the transition triplets that allow an empty machine start a new batch of any family. Let $\mathcal{A}_{2}=\set{(i,i+1,i+1):\sigma_{f}\leq i<\varphi_{f},f\in\mathcal{F}}$ be the transition triplets that allow the progression throughout the processing time of a batch of any family. Let $\mathcal{A}_{3}=\set{(\varphi_{f},0,0):f\in\mathcal{F}}$ be the transition triplets that allow a machine that completes the processing time of a batch change to the empty state. Finally, let $\mathcal{A}_{4}=\set{(\varphi_{f},\sigma_{g},\sigma_{g}):f,g\in\mathcal{F}}$ be the transition triplets that allow a machine that completes the processing time of a batch start the processing of a new batch, even of different family.
Model 2 presents the AU model using the syntax of CP-SAT [41], an open-source solver in Google OR-Tools [43]. For the variables of this model, equation (2) defines the start time of job $j\in\mathcal{J}$ on machine $m\in\mathcal{M}$. Equation (2) defines a boolean variable $p_{jm}$ that takes the value of 1 if job $j$ is processed on machine $m$ or 0 if not. Equation (2) defines the interval variable $x_{jm}$ with presence literal $p_{jm}$, which represents job $j$ being processed on machine $m$. Equation (2) defines a modified start time $\hat{s}_{jm}$ of the job $j$ on machine $m$ which is the same start time $s_{jm}$ if the job is processed on the machine (i.e., $p_{jm}=1$), or the horizon’s limit $H$ if not. Equation (2) defines the transition values of machine $m$’s automaton at each time step $t$ of the scheduling horizon $\mathcal{H}$. Finally, equation (2) defines the transition value $\hat{\tau}_{jm}$ at the modified start time $\hat{s}_{jm}$ on machine $m$, which is the initial state $\sigma_{f_{j}}$ if job $j$ is processed on machine $m$, or the empty state otherwise.
Model 2 Automaton model
Variables:
$s_{jm}$$\in\set{a\in\mathcal{H}:r_{j}\leq a},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(10)$
$p_{jm}$$\in\set{0,1},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(11)$
$x_{jm}$ $=[s_{jm},s_{jm}+{\delta}_{j})\text{ with }p_{jm},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(12)$
$\hat{s}_{jm}$ $=s_{jm}\cdot p_{jm}+H\cdot(1-p_{jm}),$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(13)$
$\tau_{mt}$ $\in\mathcal{N},$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},t\in\mathcal{H};$ $(14)$
$\hat{\tau}_{jm}$ $=\sigma_{f_{j}}\cdot p_{jm}\in\Lambda\cup\{0\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(15)$
Formulation:
minimize $\text{minimize\leavevmode\nobreak\ }\sum_{j\in\mathcal{J}}\sum_{m\in\mathcal{M}}\omega_{j}\cdot p_{jm}\cdot(s_{jm}+{\delta}_{j}),$ $(16)$
subject to,
$\forall\leavevmode\nobreak\ j\in\mathcal{J}:$ (17)
      exactlyOne ($\left(\{p_{jm}\}_{m\in\mathcal{M}}\right);$
$\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}:$ (18)
cumulative ($\left(\{x_{jm}\}_{j\in\mathcal{J}_{f}},\{{n}_{j}\}_{j\in\mathcal{J}_{f}},u_{f}\right);$
$\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M}:$ (19)
element $\left(\hat{s}_{jm},\{\tau_{mt}\}_{t\in\mathcal{H}},\hat{\tau}_{jm}\right);$
$\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}:$ (20)
automaton ($\left(\{\tau_{mt}\}_{t\in\mathcal{H}},0,\Omega\cup\{0\},\mathcal{A}\right).$
Objective function (2) minimizes the TWCT. Constraints (2) use the global constraint $\texttt{exactlyOne}(P)$, which receives a set of boolean variables $P$ and gives the value of 1 to only one of them. Hence, constraints (2) ensure that each job is processed on exactly one machine. Constraints (2) use the global constraint $\texttt{cumulative}(V,D,u)$, which receives a set of interval variables $V$, their demands $D$, and the capacity $u$ of a resource. This constraint ensures that the demands of overlapping intervals respect the resource capacity. Hence, constraints (2) ensure the maximum batch size of every family. Constraints (2) use the global constraint $\texttt{element}(t,T,\tau)$, which receives an integer decision variable $t$, a sequence of integer decision variables $T$, and an integer decision variable $\tau$. This constraint ensures that the $t$th element of the sequence $T$ takes the value of $\tau$ (i.e., $T_{t}=\tau$). Hence, constraints (2) ensure that if job $j$ is processed on the machine $m$, then the machine’s transition value at time $s_{jm}$ is the initial automaton state $\sigma_{f_{j}}$, (i.e., $\tau_{m,s_{jm}}=\sigma_{f_{j}}$); but if the job is not processed on such machine, then the transition value at the end of the scheduling horizon leaves the machine empty (i.e., $\tau_{m,H}=0$). Thus, constraints (2) guarantee that every time a job starts on a machine, the right transition value is provided to the machine’s automaton. Finally, constraints (2) use the global constraint $\texttt{automaton}(T,\sigma,\Omega,\mathcal{A})$ which receives a sequence of transition variables $T$, an initial state $\sigma$, a set of possible end states $\Omega$, and a set defining the transition triples $\mathcal{A}$ of the automaton. This constraint ensures that starting from the initial state $\sigma$, the transition variables $T$ take values that respect the automaton’s structure and end in one of the final states $\Omega$. Thence, constraints (2) ensure that the machines start empty and transition their states respecting the automaton’s structure, until reaching one of the final states $\Omega$ at the end of the scheduling horizon, or remain empty. Since there is only one possible transition after reaching the initial state $\sigma_{f}$ of any family $f$, this constraint will ensure that jobs of the same family processed on each machine don’t interrupt any other job already started. Therefore, due to the automaton’s structure, all the jobs in the same batch start and end together.
To minimize the makespan, objective function (2) should be replaced by objective function (21), which minimizes the maximum end time of the jobs.
minimize $\text{minimize\leavevmode\nobreak\ \leavevmode\nobreak\ }\max_{j\in\mathcal{J},m\in\mathcal{M}}\leavevmode\nobreak\ p_{jm}\cdot(s_{jm}+{\delta}_{j})$ (21)
Figure 3 presents the optimal solution of the AU model minimizing the TWCT. Empty states don’t have the id. The initial state of the family is in blue and the final state is in magenta. The transition values are below the arcs. Note that in each batch, both jobs send the same transition value to the automaton, which aligns them.
Figure 3: AU optimal solution minimizing TWCT
VI-B The Assign-and-Schedule Model
@@
The following three models make explicit use of the ordered set of possible batches =$\mathcal{B}=\set{1,2,3\ldots}$. The family $f^{b}\in\mathcal{F}$ of each possible batch $b\in\mathcal{B}$ is defined a priori. In addition, the model only allows jobs $\mathcal{J}_{f^{b}}$ of family $f^{b}$ to be processed on batch $b$. Hence, this set can be further partitioned into the mutually exclusive sets $\mathcal{B}_{f}=\set{b\in\mathcal{B}:f^{b}=f}\subset\mathcal{B}$ of possible batches for each family $f\in\mathcal{F}$. In this way, the first $|\mathcal{B}_{1}|$ elements of $\mathcal{B}$ correspond to the possible batches where jobs of family $1$ can be processed; the next $|\mathcal{B}_{2}|$ elements to the possible batches for jobs of the family $2$, and so on. The number of possible batches of each family is upper bounded by the number of jobs that belong to that family, hence, we set $|\mathcal{B}_{f}|=|\mathcal{J}_{f}|$.
Model 3 The Assign-and-Schedule Model Inspired by [8]
@@
$y_{b}$$\in\{[a,a+{\delta}^{f^{b}}):a\in\mathcal{H}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(23)$
$y_{bm}$$\in\{[a,a+{\delta}^{f^{b}}):a\in\mathcal{H}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(24)$
@@
$\sum_{j\in\mathcal{J}_{f^{b}}}{n}_{j}\cdot z_{jb}\nobreak\leavevmode$ $\leq u_{f^{b}}\cdot\texttt{presenceOf}\left(y_{b}\right);$ $(30)$
@@
Model 3 presents the Assign-and-Schedule (AS) model using the syntax of the commercial solver CP Optimizer. However, the same model can be implemented in open-source solvers. It is based on the CP model by Nascimento et al. [8] for AM processes, which deals with a single compatible family and does not include capacity constraints. This paper extends their formulation to handle multiple incompatible families and include the necessary constraints to ensure the maximum batch sizes. The model name originates from its clear distinction between assignment and scheduling decisions. While the job assignment to batches is handled with binary variables, the batch sequencing is handled with CP scheduling logic.
Equation (3) defines a binary variable $z_{jb}$ that takes the value of 1 if job $j\in\mathcal{J}$ is assigned to batch $b\in\mathcal{B}_{f_{j}}$, or 0 otherwise. Equation (3) defines an optional interval variable $y_{b}$ of size ${\delta}^{f^{b}}$ that represents batch $b\in\mathcal{B}$. It is optional because a batch with no jobs assigned should not be present in the solution. Equation (3) defines an optional interval variable $y_{bm}$, also of size ${\delta}^{f^{b}}$, that represents batch $b$ being processed on machine $m\in\mathcal{M}$. Equation (3) defines an interval sequence variable $\varphi_{m}$, which represents the sequence of batches scheduled on machine $m$. This sequence variable can be any permutation of the batch intervals $\set{y_{bm}}_{b\in\mathcal{B}}$ on the machine $m$ present in the solution.
Objective function (3) minimizes the TWCT. Constraints (3) assign each job to exactly one possible batch. Constraints (3) deactivate the presence of the batches if no jobs are assigned to them. Constraints (3) activate the presence of the batches if at least one job is assigned to them. Constraints (3) ensure that if a batch is used, its size respects the maximum batch capacity. Constraints (3) guarantee that batches are scheduled once all the jobs in them are available. Constraints (3) schedules each batch on exactly one machine, if present. Finally, constraints (3) use the global constraint noOverlap, which ensures that the present interval variables in the provided sequence do not overlap. Hence, constraints (3) ensure that only one batch is processed at a time on each machine.
To minimize the makespan, objective function (3) should be replaced by objective function (34), which minimizes the maximum end time of the jobs.
minimize $\text{minimize\leavevmode\nobreak\ \leavevmode\nobreak\ }\max_{j\in\mathcal{J},m\in\mathcal{M}}\leavevmode\nobreak\ p_{jm}\cdot(s_{jm}+{\delta}_{j})$ (34)
Figure 4 presents the optimal solution of the AS model minimizing the TWCT. Note that the structure of the solution changes, and there is a clear division between assignment and scheduling decisions.
@@
VI-C The Synchronized Model
Model 4 presents the Synchronized (S) model using the syntax of the commercial solver CP Optimizer. The same model can be implemented in open-source solvers with minimal changes. It is based on the CP model by Cakici et al. [10] for AM processes, which deals with a single compatible family; this paper extends their formulation to handle multiple incompatible families. Equation (4) defines an optional interval variable $x_{jm}^{b}$, of size ${\delta}_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m\in\mathcal{M}$. Equation (4) defines an optional interval variable $y_{bm}$ of size ${\delta}^{f^{b}}$ that represents the option of batch $b\in\mathcal{B}$ being processed on machine $m$. Equation (4) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (4) defines a cumulative function ${\mathscr{l}}_{m}^{b}$ that represents the load on machine $m$ due to the jobs in batch $b$.
Objective function (4) minimizes the TWCT. Constraints (4) assign each job to exactly one batch on exactly one machine. Constraints (4) allow each batch to be processed in at most one machine. Constraints (4) use the synchronize global constraint, which takes two inputs: an interval variable $v=x_{jm}^{b}$ and a set of interval variables $B=\set{y_{bm}}$. If $v$ is present, this constraint attempts to synchronize it with any intervals in $B$ that are also present [40], i.e., $\texttt{presenceOf}(v)=1\implies\texttt{startOf}(v)=\texttt{startOf}(b)\wedge\texttt{endOf}(v)=\texttt{endOf}(b),\forall\leavevmode\nobreak\ b\in B$ s.t. $\texttt{presenceOf}(b)=1$. In practice, when the interval variable representing the job being assigned to a batch on a machine (i.e., $x_{jm}^{b}$) is part of the solution, the synchronize constraint attempts to align its start and end times with those of the batch on the machine (i.e., $y_{bm}$), if it is also present. However, the synchronize constraint only ensures synchronization if both interval variables are present; it does not inherently enforce the presence of $y_{bm}$ when $x_{jm}^{b}$ is scheduled. This subtle issue, where the presence of $x_{jm}^{b}$ should logically imply the presence of $y_{bm}$ is addressed in the original model by Cakici et al. [10] by means of an additional constraints that defines the correct duration of the batches depending on the jobs assigned to it, if any. This constraint is necessary in their AM application as the duration of the batch depends on the size of the parts 3D-printed in it. However, this constraint is not required in the context of the diffusion area studied in this paper since the duration of the batches is already predefined. Nonetheless, the presence of the correct batches is still needed to be enforced. To resolve this, constraints (4) introduce an explicit implication constraint which ensures that, whenever $x_{jm}^{b}$ is included in the solution, the corresponding batch interval $y_{bm}$ is also present. Constraints (4) use the noOverlap global constraint to ensure that the present interval variables in the machine sequence do not overlap, guaranteeing at most one batch being processed at a time on each machine. Finally, constraints (4) guarantee that the load of the batches respect the maximum capacity.
Model 4 The Synchronized Model Inspired by [10]
@@ VI-C The Synchronized Model
$y_{bm}$$\in\{[a,a+{\delta}^{f^{b}}):a\in\mathcal{H}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(36)$
@@
${\mathscr{l}}_{m}^{b}$ $=\sum_{j\in\mathcal{J}_{f^{b}}}\text{pulse}\left(x^{b}_{jm},{n}_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},b\in\mathcal{B}.$ $(38)$
@@
$\forall\leavevmode\nobreak\ m\in\mathcal{M},b\in\mathcal{B}:\quad{\mathscr{l}}_{m}^{b}\leq u_{f}^{b}.$ $(45)$
To minimize the makespan, objective function (4) should be replaced by objective function (46), which minimizes the maximum end time of the jobs.
minimize $\text{minimize\leavevmode\nobreak\ \leavevmode\nobreak\ }\max_{b\in\mathcal{B},m\in\mathcal{M}}\leavevmode\nobreak\ \texttt{endOf}(y_{bm})$ (46)
Figure 5 shows the optimal solution of the S model. Note how the jobs belong now to batches and the cumulative loads are now tallied by batches.
@@
VI-D The Redundant Synchronized Model
Model 5 presents the Redundant Synchronized (RS) model using the syntax of the commercial solver CP Optimizer. The same model can be implemented in open-source solvers with minimal changes.
Model 5 Redundant Synchronized model